Which of the following pairs represent the same rational number?
 (-3/5) and (-12/20)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to check the given pair represents the same rational number.
​Then,
-3/5 = - 12/20
-3/5 = -3/5
∵-3/5 = -3/5
∴-3/5= -12/20
So, the given pair is represents the same rational number.

​​Thanks